OpenStudy (anonymous):

Solve |x + 2| = 10 {-8, 8} {-12, 8} {-12, 12}

OpenStudy (solomonzelman):

B

OpenStudy (solomonzelman):

-12 or 8

OpenStudy (austinl):

Easy peasy! (x+2)=10 -(x+2)=10 Solve for x in each of them.

OpenStudy (anonymous):

thankx all of you...